Given that 3 + i is a root of the equation z
3 − 3z
2 − 8z + 30 = 0, find the
remaining roots.
1
Expert's answer
2020-05-20T19:54:13-0400

The equation is given as:

"z^3-3z^2-8z+30=0"

Let's type the equation in the form of:

"z^3-6z^2+3z^2+10z-18z+30=0"

After some substitutions:

"z^3-6z^2+10z+3z^2-18z+30=0"

Realease "z" from the first three components, and 3 from the remaining ones:

"z(z^2-6z+10)+3(z^2-6z+10)=0"

Then, release "z^2-6z+10" from both multiplications, yields

"(z+3)(z^2-6z+10)=0"

The root of the first bracket is:

"z=-3" ;

As for the second bracket, we know that complex roots occur in conjugate pairs:

as "3+i" is a root, so must the other be "3-i"

Now, sum of the roots "=3+i+3-i=6" and,

the product of the roots "=(3+i)(3-i)=9-3i+3i-i^2=9+1=10"

Hence, the roots of the equation are: "-3; 3-i; 3+i"